Your teacher is giving a test worth 150 points. There are 46 three and five point question on the test. How many of each question aron the test?
Answer by stanbon(75887)

Your teacher is giving a test worth 150 points. There are 46 three and five point question on the test. How many of each question aron the test?
---------------
Quantity equation: T + F = 46
Point equation: 3T + 5F = 150
---------------------------------------
Multiply thru 1st equation by 3:
3T + 3F = 138
3T + 5F = 150
----------------------
Subtract 1st from 2nd and solve for "F":
2F = 12
F = 6 (# of five point questions)
--------------------------
Since T+F = 46, T= 40 (# of three point questions)
=======================
Cheers,
Stan H.